OpenStudy (anonymous):

Solve algebraically: 4x^2-9y^2=108 xy= -12 [please help] medal given.

OpenStudy (loser66):

xy =-12 (no need to set up condition that x , y not =0) ---> \[y=\frac{-12}{x}\]---> \[y^2 = \frac{144}{x^2}\] replace into the first equation to get \[4x^2-9(\frac{144}{x^2})=108\] ---> \[4x^4 - 1296 = 108 x^2\] \[4x^4 -108x^2 -1296 =0\] I dont' think it's hard for you from this

OpenStudy (anonymous):

got it! thanks @Loser66